# Inequality in Congestion Games with Learning Agents

### v1

@@ 5.2. The Amsterdam Metro Network Game
Player cost is modeled as a function of congestion using a volume–delay function widely applied in transport planning das_link_cost_function.
@@
where $t_{e}^{0}$ is the free-flow travel time and $K_{e}$ is the edge capacity. Free-flow times are calculated from real-world schedules (details in the Appendix). This function is appropriate for public transport networks, where both travel time and crowding influence modal choice li_crowding_pt. Moreover, empirical evidence from Amsterdam shows nearly half of commuters adjust behaviour in response to overcrowding monitor_sociale_2023. For simplicity, we set a uniform capacity $K_{e}=N$ across all edges, implying that the cost of an edge doubles when it is fully occupied compared to when it is empty.
Consistent with the Braess paradox, the total system cost $C$ is the sum of individual link costs (Eq. 14). Figure 2 (Top) shows the three network phases and their free-flow travel times $t_{e}^{0}$. Similarly to the Braess Network, the metro game admits at least one pure-strategy Nash equilibrium (see Appendix for detailed construction of the potential functions and full proof of equilibrium existence). Table 1 reports NE, SO, PoA, and SD, assuming no learning agents. In Figure 2 (Bottom), we show the flows in the network under the NE. A key distinction between the Amsterdam and Braess networks is that Amsterdam does not exhibit a Price of Anarchy: NE and SO coincide in all phases, and selfish routing is efficient under classical game theory assumption.
Analyzing the NE further, network extensions consistently improve fairness and efficiency. The Source Disparity decreases, and PoA remains at $1$. With the North–South line, agents from West switch to the West $\to$ South $\to$ Central route, decongesting Central–Amstel, consistent with empirical observations in Amsterdam, where North-South line became the most used one after its opening gvb_north_south_line. Phase C shifts congestion slightly back to Amstel–Central, but to a lesser extent than in Phase A. We explore whether learning agents converge to a similar equilibrium.
5.3. Experimental Setup
@@ 6. Results
6.1. Results on the Braess Network
6.1.1. Unequal learning rates amplify disparities
When agents have equal learning rates (Figure 3, top), the system converges to a state whose total cost lies between the SO and the NE. However, source disparity oscillates around zero, indicating that agents from different sources experience similar travel times. In other words, equal learning rates prevent persistent inequality.
In contrast, when learning rates differ between sources (Figure 3, bottom), disparities emerge. Agents with higher learning rates achieve lower travel times than slower-learning counterparts. This advantage persists throughout training and converges to a state of nonzero source disparity. Notably, this occurs even though the system’s overall price of learning remains similar, demonstrating that looking at aggregate efficiency metrics alone obscures underlying inequities.
@@ 6.1.1. Unequal learning rates amplify disparities
Figure 3. Top: With equal learning rates, the system converges to a profile between the Social Optimum and the Nash Equilibrium. Source disparity fluctuates around zero, indicating fairness across sources. Bottom: With unequal rates, agents with higher learning rates gain a persistent advantage. The system converges to inequitable states even though the price of learning is comparable.
Figure 4. When varying the relative learning rates between the two sources, disparities emerge after the intervention: agents with slower learning rate experience worse travel times than the faster-learning ones. This effect is most pronounced when learning rates differ sharply (e.g., 5:1 or 1:5).
6.1.2. Network interventions exacerbate disparities under unequal learning rates
Figure 5. Training under uniform learning rates in the Amsterdam Metro network. We show the evolution of social cost and source disparity over training, for Phases A, B, and C. While the system eventually converges to the Nash Eq., learning inefficiencies and oscillations in source disparity highlight the unequal effects of adaptation even under homogeneous learning.
Figure 4 shows how source disparity evolves before and after the fast intervention. To isolate the effects of heterogeneous learning rates, we fixed the average population learning rate at $\alpha=0.2$ and varied the relative rates between sources. The intervention introduces disadvantages against the slower-learning agents, resulting in higher average travel times compared to faster-learning agents. This disparity intensifies when learning rates differ sharply (e.g., 5:1 or 1:5), producing pronounced inequities.
@@ 6.1.2. Network interventions exacerbate disparities under unequal learning rates
6.2. Results on the Amsterdam Metro Network
6.2.1. Convergence under uniform learning rates exhibits learning inefficiencies but converges to the NE
@@
Figure 6. Training under heterogeneous learning rates between West and East sources in the Amsterdam Metro network. Each plot shows the evolution of price of learning and source disparity over training steps for different learning rate ratios. Unequal learning rates produce persistent inefficiencies and systematic disparities, particularly after network expansions (Phase C), demonstrating that faster-adapting commuters gain a consistent advantage over slower learners.
6.2.2. Unequal learning rates raise inefficiencies
@@
6.2.3. Persistent inefficiencies and unfairness emerge after interventions
@@ 7. Conclusion
This paper contributes to the ongoing research on fair transport network design martens2016transport; BEHBAHANI2019171; michailidis2023balancing, by emphasizing the importance of considering sources of heterogeneity that dictate mobility behaviour. We highlight the interplay between individual learning and systemic fairness. Our results suggest that interventions aimed at mitigating congestion should also account for the heterogeneity in agents’ learning, and overlooking that reinforces inequities.
Future work could extend our analysis to asymmetric networks where, even with fully rational agents, a trade-off exists between optimal and fair routing—as demonstrated in roughgarden2002unfair. Furthermore, it would be relevant to consider alternative learning algorithms, inspired in psychology and economics literature, such as Roth-Erev learning erev1998predicting or model-based algorithms daw2011model. Investigating how such inequities might be mitigated in the presence of learning agents remains an important direction for research.
{acks}
This project was funded by the Innovation Center for Artificial Intelligence (ICAI).

### v2

@@ 5.2 The Amsterdam Metro Network Game
Player cost is modeled as a function of congestion using a volume–delay function widely applied in transport planning [19].
@@
where $t_{e}^{0}$ is the free-flow travel time and $K_{e}$ is the edge capacity. Free-flow times are calculated from real-world schedules (details in the Appendix). This function is appropriate for public transport networks, where both travel time and crowding influence modal choice [30]. Moreover, empirical evidence from Amsterdam shows nearly half of commuters adjust behaviour in response to overcrowding [38]. For simplicity, we set a uniform capacity $K_{e}=N$ across all edges, implying that the cost of an edge doubles when it is fully occupied compared to when it is empty.
Consistent with the Braess paradox, the total system cost $C$ is the sum of individual link costs (Eq. 14). Figure˜2 (Top) shows the three network phases and their free-flow travel times $t_{e}^{0}$. Similarly to the Braess Network, the metro game admits at least one pure-strategy Nash equilibrium (see Appendix for detailed construction of the potential functions and full proof of equilibrium existence). Table˜1 reports NE, SO, PoA, and SD, assuming no learning agents. In Figure˜2 (Bottom), we show the flows in the network under the NE. A key distinction between the Amsterdam and Braess networks is that Amsterdam does not exhibit a Price of Anarchy: NE and SO coincide in all phases, and selfish routing is efficient under classical game theory assumption.
Analyzing the NE further, network extensions consistently improve fairness and efficiency. The Source Disparity decreases, and PoA remains at $1$. With the North–South line, agents from West switch to the West $\to$ South $\to$ Central route, decongesting Central–Amstel, consistent with empirical observations in Amsterdam, where North-South line became the most used one after its opening [27]. Phase C shifts congestion slightly back to Amstel–Central, but to a lesser extent than in Phase A. We explore whether learning agents converge to a similar equilibrium.
5.3 Experimental Setup
@@ 6 Results
6.1 Results on the Braess Network
6.1.1 Unequal learning rates amplify disparities
When agents have equal learning rates (Figure˜3, top), the system converges to a state whose total cost lies between the SO and the NE. However, source disparity oscillates around zero, indicating that agents from different sources experience similar travel times. In other words, equal learning rates prevent persistent inequality.
In contrast, when learning rates differ between sources (Figure˜3, bottom), disparities emerge. Agents with higher learning rates achieve lower travel times than slower-learning counterparts. This advantage persists throughout training and converges to a state of nonzero source disparity. Notably, this occurs even though the system’s overall price of learning remains similar, demonstrating that looking at aggregate efficiency metrics alone obscures underlying inequities.
@@ 6.1.1 Unequal learning rates amplify disparities
Figure 3: Top: With equal learning rates, the system converges to a profile between the Social Optimum and the Nash Equilibrium. Source disparity fluctuates around zero, indicating fairness across sources. Bottom: With unequal rates, agents with higher learning rates gain a persistent advantage. The system converges to inequitable states even though the price of learning is comparable.
Figure 4: When varying the relative learning rates between the two sources, disparities emerge after the intervention: agents with slower learning rate experience worse travel times than the faster-learning ones. This effect is most pronounced when learning rates differ sharply (e.g., 5:1 or 1:5).
6.1.2 Network interventions exacerbate disparities under unequal learning rates
Figure 5: Training under uniform learning rates in the Amsterdam Metro network. We show the evolution of social cost and source disparity over training, for Phases A, B, and C. While the system eventually converges to the Nash Eq., learning inefficiencies and oscillations in source disparity highlight the unequal effects of adaptation even under homogeneous learning.
Figure˜4 shows how source disparity evolves before and after the fast intervention. To isolate the effects of heterogeneous learning rates, we fixed the average population learning rate at $\alpha=0.2$ and varied the relative rates between sources. The intervention introduces disadvantages against the slower-learning agents, resulting in higher average travel times compared to faster-learning agents. This disparity intensifies when learning rates differ sharply (e.g., 5:1 or 1:5), producing pronounced inequities.
@@ 6.1.2 Network interventions exacerbate disparities under unequal learning rates
6.2 Results on the Amsterdam Metro Network
6.2.1 Convergence under uniform learning rates exhibits learning inefficiencies but converges to the NE
@@
Figure 6: Training under heterogeneous learning rates between West and East sources in the Amsterdam Metro network. Each plot shows the evolution of price of learning and source disparity over training steps for different learning rate ratios. Unequal learning rates produce persistent inefficiencies and systematic disparities, particularly after network expansions (Phase C), demonstrating that faster-adapting commuters gain a consistent advantage over slower learners.
6.2.2 Unequal learning rates raise inefficiencies
@@
6.2.3 Persistent inefficiencies and unfairness emerge after interventions
@@ 7 Conclusion
This paper contributes to the ongoing research on fair transport network design [31, 4, 32], by emphasizing the importance of considering sources of heterogeneity that dictate mobility behaviour. We highlight the interplay between individual learning and systemic fairness. Our results suggest that interventions aimed at mitigating congestion should also account for the heterogeneity in agents’ learning, and overlooking that reinforces inequities.
Future work could extend our analysis to asymmetric networks where, even with fully rational agents, a trade-off exists between optimal and fair routing—as demonstrated in [42]. Furthermore, it would be relevant to consider alternative learning algorithms, inspired in psychology and economics literature, such as Roth-Erev learning [23] or model-based algorithms [20]. Investigating how such inequities might be mitigated in the presence of learning agents remains an important direction for research.
Acknowledgments
This project was funded by the Innovation Center for Artificial Intelligence (ICAI) and the City of Amsterdam.